### v1

# QUIS: Question-guided Insights Generation for Automated Exploratory Data Analysis

Abstract
Discovering meaningful insights from a large dataset, known as Exploratory Data Analysis (EDA), is a challenging task that requires thorough exploration and analysis of the data. Automated Data Exploration (ADE) systems use goal-oriented methods with Large Language Models and Reinforcement Learning towards full automation. However, these methods require human involvement to anticipate goals that may limit insight extraction, while fully automated systems demand significant computational resources and retraining for new datasets. We introduce QUIS, a fully automated EDA system that operates in two stages: insight generation (ISGen) driven by question generation (QUGen). The QUGen module generates questions in iterations, refining them from previous iterations to enhance coverage without human intervention or manually curated examples. The ISGen module analyzes data to produce multiple relevant insights in response to each question, requiring no prior training and enabling QUIS to adapt to new datasets.
QUIS: Question-guided Insights Generation for Automated Exploratory Data Analysis
Abhijit Manatkar 1, Ashlesha Akella1,2, Parthivi Gupta1, Krishnasuri Narayanam1, 1IBM Research, India, 2Indian Institute of Technology Kharagpur, India,
1 Introduction
Exploratory Data Analysis (EDA) is the process of discovering meaningful insights from vast amounts of data, and it is a complex task requiring careful data exploration. There are various EDA techniques to uncover insights by analyzing patterns in the data. Automated Data Exploration (ADE) systems accelerate the EDA process through automation.
ADE literature includes statistics-based Sellam et al. (2015); Ding et al. (2019); Wang et al. (2020); Ma et al. (2021, 2023) and interactive methods Milo and Somech (2016, 2018b); Agarwal et al. (2023); He et al. (2024), where users explore data through natural language queries or receive suggestions for subsequent actions. Visualization-based techniques Vartak et al. (2015); Demiralp et al. (2017); Srinivasan et al. (2018); Wu et al. (2024) offer visual insights and allow further queries. However, these methods can become resource-intensive due to extensive user interactions. Goal-oriented ADE approaches, generate insights based on predefined objectives Tang et al. (2017); Seleznova et al. (2020); Omidvar-Tehrani et al. (2022); Laradji et al. (2023). This approach directs the exploration using predefined objectives, such as natural language goals or statistical measures of interestingness. While this reduces user interactions, it may constrain the insights to only those aligned with the predetermined goals.
ADE using reinforcement learning is studied Milo and Somech (2018a); Bar El et al. (2019, 2020); Personnaz et al. (2021); Garg et al. (2023) to achieve full automation. While these systems minimize user involvement, they often demand dataset-specific training and substantial computational resources, particularly as the number of features, categorical values, or patterns increases, making the process increasingly challenging.
1.1 Motivation
An effective EDA system exercises statistical examination with attention to data semantics, such as analyzing trends in date and sales price or examining the impact of weather on flight delay. Systems like Demiralp et al. (2017); Deutch et al. (2022); Ma et al. (2023); Guo et al. (2024) leverage Large Language Models (LLMs) to drive the analysis based on natural language goals. Systems which use LLMs to generate relevant questions based on natural language goals Laradji et al. (2023), drive insight discovery based on user queries Wang et al. (2022), and interpret analysis objectives from the user’s natural language input to specify desired outcomes Lipman et al. (2024) have also been proposed. Guiding EDA through insightful questions enables purposeful exploration, clarifying analysis goals, and deriving actionable insights. In contrast, such a goal-oriented approach Laradji et al. (2023) may overlook unanticipated critical findings.
1.2 Our Contributions
We propose a two-stage ADE system, QUIS, that fully automates the EDA process. In the first stage, QUIS generates questions based solely on the data semantics (dataset information like name, description, column names, and column descriptions) without requiring predefined objectives. In the second stage, QUIS uses statistical analysis to produce insights corresponding to the questions from the first stage. This research contributes to the following advancements
  • Question Generation (QUGen) module generates questions in iterations, where questions generated in previous iterations, along with their reasoning and relevant information, serve as examples for subsequent iterations. This approach helps generate unique questions with broader coverage by providing additional context and guidance to the LLM in each iteration. Our approach eliminates the dependency on manually curated examples and predefined analysis goals.
  • Insight Generation (ISGen) module analyzes the data using statistical patterns and classical search techniques to generate insights in response to the questions from the QUGen module without requiring prior training. For a given question, this module provides multiple relevant insights.
QUIS offers notable benefits, including reduced dependency on expert knowledge, enhanced efficiency in the exploration process, the ability to uncover a broader range of insights from the data, and ease of use across various datasets.
Figure 1: The Question Generation (QUGen) module of QUIS system generates questions refined over iterations using data semantics, while the Insight Generation (ISGen) module generates insights (bottom-right) using those questions via statistical analysis. Question is encapsulated inside the Insight Card.
2 Preliminaries
Although it is challenging to precisely define the notion of an insight due to variations in users’ objectives, for this work, we adopt the definition of an insight consistent with previous studies Ding et al. (2019); Ma et al. (2023). Consider a tabular dataset $D=\{X_{1},X_{2},\dots X_{n}\}$ where each $X_{i}$ is an attribute (column) of the dataset. An insight, denoted by $Insight(B,M,S,P)$, consists of the following:
  1. 1.
    Perspective - A perspective consists of a tuple $(B,M)$. $B$ represents the breakdown attribute, and $M$ is the measure, referring to a quantity of interest from the table. Typically, $M$ is of the form $agg(C)$ where $agg$ (measure function) is an aggregation function , like count(), mean(), sum(), etc., and $C$ (measure column) is a numerical attribute of the dataset. $B$ is the breakdown dimension, a column of interest from the table, for which we want to compare different values of $M$. For each perspective $(B,M)$, we can compute a view $view(D,B,M)$ of the dataset $D$ by grouping on $B$ and calculating the measure $M$ for each group. For example, computing $view(D,\texttt{Year},\texttt{mean}(\texttt{Performance}))$ is equivalent to applying the SQL query: SELECT Year, AVG(Performance) FROM $D$ GROUP BY Year.
  2. 2.
    Subspace - A subspace $S=\bigcup_{i}\{(X_{i},y_{ik})\}$ is a set of filters that determine a subset ($D_{S}$) of the dataset D. Each $X_{i}$ is an attribute, and each $y_{ik}$ is a corresponding value of the column $X_{i}$ of $D$. A tuple $(X_{i},y_{ik})$ denotes that the dataset is to be filtered for rows where $D[X_{i}]=y_{ik}$.
  3. 3.
    Pattern - The pattern $P$ represents the type of insight observed. It belongs to a predefined set of known patterns, such as trends or outliers.
The QUIS system incorporates the following insight types as candidates for our patterns: (1) Trend - An increasing or decreasing trend is seen in a set of values. (2) Outstanding Value - The largest (or smallest) value in a set of values is significantly larger (or smaller) than all other values in the set. (3) Attribution - The highest value accounts for a large proportion ($\geq 50\%$) of the total of all values in the set. (4) Distribution Difference - The distribution of values in a set changes notably from one subspace to another.
As an example, consider the insight given by
  • $B=\texttt{Year}$, $M=\texttt{mean}(\texttt{Performance})$
[…]
  • $P=\texttt{Trend}$
This insight suggests that for the "Sales" department, there has been a trend in the average employee performance over the years.
By combining a breakdown $B$, a measure $M$, and a subspace $S$, we can compute a unique view of the dataset $D$ by first applying the filters in $S$ on $D$ to arrive at $D_{S}$, then computing the $view(D_{S},B,M)$ as described. Let ${\mathbb{V}}(D)$ be the set of all possible views of dataset $D$ that can be computed in this manner. A search for insights involves finding views belonging to ${\mathbb{V}}(D)$ for which an insight pattern $P$ is observed. As the size of ${\mathbb{V}}(D)$ grows exponentially with the number of columns in $D$, searching for insights by enumerating all possible views in ${\mathbb{V}}(D)$ is inefficient. Therefore, it becomes important to limit the search to subspaces that are semantically meaningful and statistically relevant.
3 Methods
The EDA process is often guided by the questions that arise from the semantic context and the statistical properties of the dataset. Hence, we propose an approach, QUIS (QUestion-guided InSight generation), that employs a two-stage process (refer to Figure 1). The first stage, QUGen, leverages LLMs to formulate questions based on the dataset schema, basic statistics, and iteratively updated in-context examples. The second stage, question-driven insight generation (ISGen), systematically analyzes the tabular data statistics based on the questions to uncover meaningful insights.
3.1 Question Generation (QUGen)
Our QUIS framework begins with QUGen producing a set of Insight Cards. Each Insight Card encapsulates relevant information aligning with recent advances in automated EDA Ding et al. (2019); Ma et al. (2021). In particular, an Insight Card (example in Figure 2) includes four components: Question, which is the generated natural language question aimed at guiding data analysis; Reason, which explains the rationale behind the generated question to help further analysis; Breakdown $B$, and Measure $M$. The Reason is used by QUGen to enhance the coverage, and other components are used by both QUGen and ISGen.
Figure 2: Example Insight Card
3.1.1 Input Prompt
The prompt for QUGen consists of several key components (for details refer to Figure 6 in Appendix), starting with a high-level description of the data analysis task objective. It then provides detailed instructions for generating an Insight Card by examining the table schema and basic statistics along with a few-shot example table schemas and their sample Insight Cards. Additionally, the prompt includes the schema of the test table and concise natural language descriptions of key statistics summarizing essential information. These statistics are generated by prompting an LLM (for prompt refer Figure 7 in Appendix) with few-shot examples to generate basic statistical questions, which are transformed into SQL, applied to the dataset, and translated into natural language responses.
3.1.2 QUGen pipeline
The QUGen LLM is prompted to generate multiple Insight Cards, as shown in Figure 1. The LLM’s response is sampled $s$ times with a temperature $t$, with each sample containing $n$ Insight Cards. However, the exact number of Insight Cards per sample may vary slightly due to the fixed output token length.
Each Insight Card undergoes a filtering process: first, cards with questions not semantically relevant to the table schema are removed using semantic similarity computed using the all-MiniLM-L6-v2 Sentence Transformers model Reimers and Gurevych (2019). Next, duplicate Insight Cards are eliminated based on semantic similarity between pairs of questions. Simple or rudimentary questions are filtered out by converting them to SQL queries and applying them on the dataset; if a query returns only one row, the question is discarded. This ensures that only in-depth questions are retained for comprehensive data analysis.
QUGen is iterative in nature (refer Figure 1). It uses subset of Insight Cards generated until the current iteration as in-context examples in the prompt for the next iteration, offering supplementary context and guidance to ensure generation of unique Insight Cards distinct from that of previous iterations. A key advantage of this comprehensive approach by QUGen module is that it eliminates the need for manually providing dataset specific in-context examples, as the Insight Cards generated by the earlier iterations help the LLM understand the dataset context during the subsequent iterations. A collection of Insight Cards accumulated over a certain number (e.g., 10) of iterations are provided as the output by QUGen process.
3.2 Insight Generation (ISGen)
This module uses classical search techniques and insight scores based on different statistical measures to identify interesting insights from the data.
To determine whether a combination of $B$, $M$, and $S$ reveals a particular pattern $P$, the module uses scoring functions based on data statistics and applies appropriate thresholds. For each insight pattern $P$, a corresponding scoring function $\textsc{ScoreFunc}_{P}:{\mathbb{V}}(D)\rightarrow\mathbb{R}$ is defined, along with a threshold value $T_{P}$. Further details about the scoring function and thresholds for each pattern are provided in Appendix A. If a combination of $B$, $M$, and $S$ results in a view $v=view(D_{S},B,M)$ such that $\textsc{ScoreFunc}_{P}(v)>T_{P}$, the insight pattern $P$ is considered to have been observed in $v$.
An Insight Card produced by QUGen module is processed in two stages; first via identifying a basic insight followed by a subspace search for deeper insights as described below.
3.2.1 Basic Insight
Extraction of a basic insight helps to depict any meaningful patterns in the relationship between $B$ and $M$ considering the entire dataset without applying any filters. The basic insight is derived from an Insight Card by computing the view $v_{0}=view(D,B,M)$. The applicable insight patterns are determined based on the data type of the breakdown $B$ and the measure $M$. For instance, if $B$ is an ordinal column like Year or Revenue, then the Trend pattern becomes relevant. Then, scores corresponding to these insight patterns are evaluated. For an insight pattern $P$, if $\textsc{ScoreFunc}_{P}(v_{0})>T_{P}$, then $Insight(B,M,\phi,P)$ is returned as a basic insight (here $\phi$ is an empty set).
Algorithm 1 Insightful Subspace Search
1:Dataset $D$, Initial subspace $S_{0}$, perspective $(B,M)$, language model $LLM$, ScoreFunc, beam_width, max_depth, exp_factor
2:Top-K subspaces by score $\{S_{1},\dots S_{k}\}$
3:function expand($S$)
[…]
7:     $y\leftarrow\texttt{sample}(D[X])$
8:     return $S+(X,y)$
9:end function
10:$\texttt{beam}\leftarrow[(S_{0},\textsc{ScoreFunc}(S_{0}))]$
[…]
12:     for $(S,\texttt{score})\in\texttt{beam}$ do
[…]
14:              $S_{new}\leftarrow\textsc{expand}(S)$
15:              $\texttt{score}\leftarrow\textsc{ScoreFunc}(S_{new})$
16:              $\texttt{beam}.\texttt{add}((S_{new},\texttt{score}))$
17:         end for
18:     end for
[…]
20:end for
21:return beam
3.2.2 Subspace Search for Deeper Insights
Further insights can be generated from an Insight Card by searching for subspaces where the insight patterns are observed. To do so, we carry out a beam search procedure Russell and Norvig (2010) as described in Algorithm 1. The search takes an initial subspace $S_{0}$, a perspective $(B,M)$ and a score function $\textsc{ScoreFunc}_{P}$ corresponding to insight pattern $P$ as input. A beam of the current best subspaces is maintained. At each step, each subspace $S$ in the beam is expanded to exp_factor number of subspaces. Each expanded subspace $S_{new}$ is obtained by adding a filter $(X,y)$ to $S$. The selection of $(X,y)$ happens in two steps; selecting the filter column $X$ followed by $y$, the value to filter.
First, an LLM is prompted with $(B,M)$ and an instruction to return candidate filter columns ${\mathbb{X}}^{LLM}=\{{X^{LLM}_{1}\dots X^{LLM}_{k}}\}$ that can lead to semantically meaningful insights. $X$ is obtained by sampling from a distribution of available columns (columns of $D$ that have not been used in filters in $S$) with the candidate filter columns ${\mathbb{X}}^{LLM}$ having a probability mass of $w_{LLM}\in[0,1]$ distributed evenly over available columns with the rest of the mass ($1-w_{LLM}$) distributed over the remaining columns ($D\setminus{\mathbb{X}}^{LLM}$). $w_{LLM}$ is decided in such a way to ensure that semantically relevant columns are picked with a high likelihood for filtering while ensuring that other columns also have a chance of being picked.
After picking $X$, we need to pick a value $y$ from $D[X]$. To encourage the selection of values with higher frequency, $y$ is sampled from a distribution over the unique values $\{y_{1},\dots y_{k}\}$ in $D[X]$ where the probability $P(y_{i})$ of selecting $y_{i}$ is given by:
$P(y_{i})=\frac{\log(1+N(y_{i}))}{\sum_{i}{\log(1+N(y_{i}))}}$
$N(y_{i})$ is the frequency $y_{i}$’s appearance in $D[X]$.
Each candidate filter $S_{new}$ is evaluated by calculating $\textsc{ScoreFunc}_{P}(view(D_{S_{new}},B,M))$ (referred to as $\textsc{ScoreFunc}(S_{new})$ in Algorithm 1 for conciseness). After a round of expansion and evaluation, the beam is truncated to the top-k (subspace, score) pairs ranked by the score. This process repeats until the maximum desired depth of subspaces, then the final list of subspaces is returned.
The subspaces found in the search procedure are further filtered to only those $S$ for which $\textsc{ScoreFunc}_{P}(view(D_{S},B,M))>T_{P}$ to output an insight $Insight(B,M,S,P)$.
3.2.3 Post Processing
The post-processing stage of an insight formulates the final insight response, which consists of a natural language description and a corresponding data visualization, as shown in Figure 1 (ISGen). These components are based on the identified pattern $P$. For each pattern $P$, the natural language response uses a predefined template to clearly communicate the key findings. For details in the plotting conditions for each pattern refer to Appendix B.
4 Experimental Evaluation
In our study, we evaluated the QUIS pipeline’s effectiveness using human assessment and insight scores on three datasets: Sales Verma (2024), Adidas Sale Chaudhari (2022) and Employee Attrition Subhash (2017). Human evaluation focused on the individual insights assessing Relevance, Comprehensibility, and Informativeness (details in Appendix C). We tested two conditions: 1. OnlyStats, replacing the QUGen module with a purely statistics based card generation module, to assess the autonomous performance of ISGen, 2. QUIS, where both QUGen and ISGen were involved. For further information about the parameters of the experimental conditions, please refer to Appendix D. The insights were evaluated by six participants who are well-versed in data analysis, with each insight assessed by three different evaluators. Each criterion - relevance, comprehensibility, and informativeness - was rated on a scale of 1 to 5; where 1 indicated the insight was not relevant, comprehensible, or informative; and 5 indicated the insight was highly relevant, comprehensible, or informative.
4.1 Human Evaluation
The results of the human evaluation in Figure 3 shows that for the Sales and Employee Attrition datasets, QUIS outperformed the OnlyStats baseline in terms of relevance, comprehensibility, and informativeness, suggesting QUIS’s overall effectiveness. However, in the Adidas Sales dataset, OnlyStats performed slightly better, likely due to specific characteristics of this dataset which favour a simpler analytical approach.
Figure 3: Comparison of Average Human Evaluation Scores for QUIS and OnlyStats across 3 datasets.
4.2 Insight Score
We compare the average normalized outputs (in the range $[0,1]$) of ScoreFunc for all insights returned by the two experimental conditions. The comparison of scores across datasets shows that QUIS consistently outperformed the OnlyStats condition, with higher scores across all datasets as shown in Figure 4.
Figure 4: Comparison of Insight score for QUIS and OnlyStats.
4.3 Diverse Insight Cards
To assess the effect of the iterative process of QUIS on Insight Card diversity, we analyzed the number of unique cards generated by QUIS over multiple generations (with varied number of total iterations). We started with 1 iteration and a sampling rate of 20, then progressed to 11 iterations with a sampling rate of 2, keeping the total number of outputs generated by the LLM constant at 20. In the first condition, no few-shot examples were used, while in the last condition, QUGen iterated 10 times, appending the prompt with new few-shot examples sampled from all previous iterations (refer Figure 5).
The iterative process produced more diverse Insight Cards, as shown by the rise in the number of unique cards across successive iterations.
Figure 5: Total number of unique insight cards generated by QUIS under non-iterative (1 iteration) and iterative (up to 11 iterations).
5 Conclusion & Future Work
EDA systems often rely on user-generated, goal-oriented questions, which means the quality of the generated insights depends solely on these input questions, introducing potential overhead. To address this limitation, we propose a fully automated EDA system that generates dataset-specific questions automatically and performs insight discovery. This system operates in a data-agnostic manner, requiring no prior training, thereby minimizing the dependency on user input and streamlining the overall insight discovery process.
As a future work, we propose to enhance QUGen to generate questions in chunks where ISGen processes each chunk of questions before QUGen generates the next chunk. This would enable QUGen to use insights and their scores from previous chunks to inform the generation of subsequent chunks. Additionally, we will explore incorporating other types of insights as future work. For example, we aim to include outlier in time-series, anomaly detection, predictive insights and trend reversal to further enhance the variety and depth of insights generated by the QUIS system.
Appendix A Scoring Functions for ISGen
Let $\{v_{1},\dots v_{k}\}$ be the values for view $v\in{\mathbb{V}}(D)$ for some dataset $D$.The following scoring functions are defined for measuring the degree to which a particular pattern is seen in $v$.
  1. 1.
    Trend - The trend pattern is observed when a sequence of values either increases or decreases monotonically. For quantifying the degree to which the trend pattern is seen, we use the Mann-Kendall Trend Test Mann (1945); Kendall (1975). Specifically we use the implementation in the pyMannKendall package Hussain and Mahmud (2019). Let $MK(v)$ return the p-value calculated using the Mann-Kendall test for a $v\in{\mathbb{V}}(D)$. Then the score function is given by:
    $\textsc{ScoreFunc}_{\texttt{Trend}}(v)=1-MK(v)$
    The threshold $T_{\texttt{Trend}}$ is set to $0.95$ so that only views having a p-value $<0.05$ are returned.
  2. 2.
    Outstanding Value - The outstanding value pattern is observed when the largest (or most negative) value is much larger (or more negative) than other values. For this pattern, the scoring function calculates the ratio between the largest value in the set and the second largest value in the set. Let $v_{max_{1}}$ and $v_{max_{2}}$ be the two largest (absolute) values in the set. The score is then defined as:
    $\textsc{ScoreFunc}_{\texttt{OV}}(v)=\frac{v_{max_{1}}}{v_{max_{2}}}$
    The threshold for this pattern is set to $T_{\texttt{OV}}=1.4$
  3. 3.
    Attribution - The attribution pattern is observed when the top-value in a set of values accounts for more than 50% of the sum of all values. The score function used for this insight uses the ratio of the largest value to the sum of all values.
    $\textsc{ScoreFunc}_{\texttt{Attr}}(v)=\frac{\max(\{v_{1},\dots v_{k}\})}{\sum_{i}{v_{i}}}$
    As this pattern holds when the highest value is more than 50% of the total, the threshld is set as $T_{\texttt{Attr}}=0.5$.
  4. 4.
    Distribution Difference - This insight pattern can only be observed when the aggregation in the measure is COUNT(). Let $v^{I}$ and $v^{F}$ be the initial and final views. We use the Jensen-Shannon divergence Lin (1991) to compare the difference between the two distributions:
    $\textsc{ScoreFunc}_{\texttt{DD}}(v^{I},v^{F})=JSD(\frac{v^{I}}{\sum_{i}{v^{I}_{i}}}||\frac{v^{F}}{\sum_{i}{v^{F}_{i}}})$
    The threshold is set to $T_{\texttt{DD}}=0.2$.
Appendix B Plotting per Pattern
  • Trend: Scatter plots with trend lines are used to describe the increasing or decreasing nature of the data.
  • Outstanding Value: Bar charts are used for depicting the difference in the factors.
  • Attribution: Bar charts are used to show the percentage contribution of different factors
  • Distribution Difference: Pie charts are used to compare the distributions before and after a condition.
Appendix C Human Evaluation Criteria
The participants in our user study were asked to rate each generated insight on the following criteria on a scale of 1-5.
  • Relevance: To what extent the insight is applicable and useful in a given context?
  • Comprehensibility: To what extent is this insight understandable and easy to follow?
  • Informativeness: Does the insight provide substantial information for understanding the data?
Appendix D Experimental Conditions
D.1 OnlyStats
The OnlyStats experimental condition replaces QUGen with a purely statistical method for generating $(B,M)$ pairs as follows. First, a random $B$ is sampled from the list of all eligible columns of the table. This is followed by computing the Kruskal-Wallis test Kruskal and Wallis (1952) of association between breakdown $B$ and all possible measures $M$ in the table. The Kruskal-Wallis test is a non-parametric variance analysis test, used to determine if two sets of samples come from different distributions. The top 20 pairs of $(B,M)$, ranked according to the strength of association measured by the Kruskal-Wallis test are selected as input to ISGen.
D.2 QUIS
For QUIS, the following parameter values were used:
QUGen
  • LLM: Llama-3-70b-instruct AI@Meta (2024)
  • Sampling temperature $t=1.1$
  • Number of samples at each iteration $s=3$
  • Number of iterations $n=10$
  • Number of in-context examples = 6
ISGen
[…]
  • $w_{LLM}=0.5$
Sales Dataset
Schema Sample Questions Insights
Sales ( Retailer CHAR Region CHAR SalesMethod CHAR Product CHAR PricePerUnit INT UnitsSold INT TotalSales INT OperatingProfit INT OperatingMargin DOUBLE ) Do products with higher unit prices result in higher total revenue? [Uncaptioned image]
What is the average pricing strategy employed for each product category? [Uncaptioned image]
Table 1: QUIS Results for Sales Dataset
Employee Attrition Dataset
Schema Sample Questions Insights
Employee Attrition ( Age INT Attrition CHAR BusinessTravel CHAR DailyRate INT Department CHAR DistanceFromHome INT Education INT EducationField CHAR EmployeeCount INT EmployeeNumber INT EnvironmentSatisfaction INT Gender CHAR HourlyRate INT JobInvolvement INT JobLevel INT JobRole CHAR JobSatisfaction INT MaritalStatus CHAR MonthlyIncome INT MonthlyRate INT NumCompaniesWorked INT Over18 CHAR OverTime CHAR PercentSalaryHike INT PerformanceRating INT RelationshipSatisfaction INT StandardHours INT ) What is the relationship between employees’ Education levels and their Attrition rates? [Uncaptioned image]
What is the distribution of Attrition rates across different Departments? [Uncaptioned image]
Table 2: QUIS Results for Employee Attrition Dataset
Figure 6: QUGen Prompt Template
Figure 7: Natural Language Statistics Prompt Template.